### v6

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a quantum-inspired metaheuristic designed to address general optimization problems characterized by rugged landscapes and high-dimensionality with an efficient convergence rate. The study presents a comprehensive comparative evaluation of HEO’s performance against extensively-used optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating general optimization problems and providing valuable insights into its performance. The test of HEO in Pressure Vessel Design and Tubular Column Design infers its feasibility and potential in real-time applications. Further validation in Osmancik-97 and Cammeo Rice Classification proves the effectiveness of HEO and achieves a higher accuracy record.
@@ 1 Introduction
The main motivation behind the development of the Halfway Escape Optimization (HEO) algorithm stems from the limitations of existing optimization methods in terms of efficiency and adaptability to various industrial single-objective optimization problems. HEO is a quantum-inspired metaheuristic designed to tackle the distinction between different search spaces in industrial scenarios with a focus on achieving fast convergence rates and reducing the time cost in searching. Inspired by the behavior of quantum particles and the concept of halfway escape, the algorithm offers a versatile and adaptive approach to exploration and exploitation in challenging optimization domains.
This paper provides a comprehensive analysis and evaluation of the HEO algorithm, focusing on its core principles, adaptability, and robustness in navigating diverse optimization landscapes. The unique energy-driven behavior, vibration strategies, and exploratory mechanisms of HEO are examined in the context of solving a range of simple as well as complex benchmark functions. The evaluations demonstrate HEO’s effectiveness in achieving convergence, balancing exploration and exploitation, and discovering high-quality solutions across a variety of challenging optimization problems.
@@ 2 Related Work
There still have other meta-heuristic algorithms, those algorithms could classified as two category: bio-inspired and the nature-inspired, nature-inspired means the optimization algorithms are inspired from physical or chemical laws in nature. Bio-inspired optimizations algorithms including Whale Optimization Algorithm(WOA)[23], Artificial Bee Colony(ABC)[19], Bat Algorithm(BA)[42], Cuckoo Search(CS)[44],Dragonfly Algorithm (DA)[22], Flower Pollination Algorithm(FPA)[43], Cheetah Optimizer (CO)[3], Genetic Programming (GP)[40], Fox Optimizer(FOX)[26], GOOSE algorithm[15]. Nature-inspired optimization algorithms contains Snow Ablation Optimizer (SAO)[8],Light Spectrum Optimization (LSO)[2], Nuclear Reaction Optimization (NRO)[39],Water Cycle Algorithm (WCA)[13],Chernobyl Disaster Optimizer (CDO)[32]. Those bio-inspired as well as nature-inspired algorithms both have contributed significantly to the field of optimization, each with its unique strengths.
@@ 3 Proposed Methodology
In recent years, swarm optimization algorithms have been widely applied in various fields, such as machine learning, engineering design. These algorithms have achieved remarkable success in solving optimization problems and have shown great potential in handling complex tasks. However, these algorithms often face difficulties when faced with complex landscapes. One example of a group optimization algorithm facing a rugged landscape is the Particle Swarm Optimization (PSO) algorithm. When facing rugged landscapes with many local optima, PSO often gets trapped in these local optima and fails to find the global optimum[37].
@@
Yet QPSO and HEO are both quantum-inspired algorithms with the same mechanism in computing global weight for updating position(seen in equation (6)), they still have significant differences. In contrast to the Quantum Particle Swarm Optimization (QPSO), the Halfway Escape Optimization (HEO) algorithm aims to provide such an approach by leveraging insights from quantum patterns and halfway escape strategies.
While QPSO only allows for a single state change, HEO is more inclined to escape when convergence is achieved, resulting in the inclusion of velocity during the escape phase. Moreover, HEO incorporates random searches not only within the range of updates but also for the entire swarm, enhancing the swarm’s ability to solve unconstrained optimization problems. Additionally, utilizing a center clipping strategy in HEO contributes to its superior convergence rate. By merging these elements, HEO strives to effectively uncover high-quality solutions in a wide range of applications.
The Position Update mechanism, which forms the basis of all swarm optimization algorithms, which are the moving way of each individual entity in the swarm, is designed differently in HEO. Instead of focusing on a single peak, the quantum in HEO has two states that determine whether it should approximate the optimum. This unique way of updating positions allows the swarm in HEO to escape from false peaks, thus addressing the first difficulty mentioned earlier. Additionally, the movement of the quantum in HEO is not influenced by gradients like in the PSO algorithm, making it more stable and less likely to be hindered by complex landscapes. Furthermore, in order to navigate complex hollow landscapes, HEO incorporates mechanisms such as increasing the escape factor, Vibration, and Random Skip to accelerate the quantum’s ability to overcome obstacles. Lastly, for exploring complex situations, the Center Clipping mechanism in HEO ensures that the swarm is relatively concentrated around the already discovered optima, thereby efficiently searching for solutions in a convex environment.
@@ 3.2 Position Update
$x_{i+1}=x_{i}+v_{g}+v_{l},\quad i=1,2,3...N$ (1)
$v_{g}=(x_{g}-x_{i}(c_{i}+1)\times r_{1})\times r_{2}\times r_{3}$ (2)
$v_{l}=(x_{l}-x_{i}(c_{i}+1)\times r_{1})\times r_{2}(1-r_{3})$ (3)
@@ 3.3 Vibration
Another behavior of HEO is common in other algorithms, including AFSA, when the search is meaningless to make a random step for an individual search. Yet those algorithms adopt the strategy, they ignore the influence of size of steps, and those search approaches highly rely on the hyper-parameters tunning. HEO tries to use the standard deviation of positions to judge the magnitude of one random step, creating a relatively stable performance compared to the others, this random search is called Vibration.
$x_{i+1}=x_{i}+\frac{n}{1+e^{a_{i}}}$ (8)
@@
$n\sim N(0,\sigma_{x_{i}})$ (11)
Just like the behavior of quantum in physics, with higher energy level $a_{i}$, the quantum would be harder get influenced by others[34], which means smaller oscillations(seen in equation (8) and Fig.3), maintain an extent of diversity in the population. To maintain the convergence of the whole group, the increment of energy level $a_{i}$ only works in a few quantums. Additionally, other scaling approaches to adapting the space of high dimensions with gamma functions are computed much slower than just simply calculating the standard deviation of the position vectors. The reason for using Normal distribution due its universal applicability. According to the central limit theorem, with the increasing of the dimension, the size of the vibration is more approximate to the size of the step in the real situation. Since the HEO does not include periodical functions in this behavior, it cannot be defined as oscillations, that’s the reason for its name.
@@ 3.4 Center Clipping
As Fig.4 illustrated, the quantum not only would be restricted by the search area or constraints $S_{bound}$(the area with (a)) with half of the side length $b$, but the group also would be restricted by $S_{group}$. $S_{group}$(the area of (b) in Fig.3)is the search space of the group that HEO allows to search with half of the side length $b_{g}$, the search space in HEO are all in the shape of a square, making it computing faster than calculating sphere. Moreover, as equation (13) shows, this bound of group $S_{group}$ depends on the differences between the current quantum’s position and the best one that HEO found. Additionally, this center clipping mechanism uses a random factor $r_{5}$, which makes this bound soft edge not converge too early and maintains the diversity of the group. The final result of the group gets into the range of $S_{result}$, which is the overlap area of the $S_{group}$ and $S_{bound}$.
Figure 4: Center Clipping
@@ 3.5 Random Skip
$x_{i}=\frac{x_{i}+\vec{r}}{2}$ (15)
$\vec{r}_{j}\in\vec{r}$ (16)
$\vec{r}_{j}\sim U(0,b)$ (17)
As Fig.5 shows, the random skip behavior in HEO allows the entire group of quantum to explore a completely new and random search space, the blue area represents the search space, and the orange part is for the constraint. This is achieved by randomly shifting the position of each quantum by a random vector $\vec{r}$ in Equation (15). The components of the random vector $\vec{r}_{j}$ are uniformly sampled from the interval $[0,b]$, as shown in equation (17). There are numerous $b$ for $\vec{r}$, hence the shape of the search space also changed with it as in sub-figure(a)(b) does, also the space restricts the constraints as sub-figure(c) is shown. This random skip mechanism enables HEO to explore different regions of the search space and potentially discover new and better solutions.
Figure 5: Random Skip
@@ 5.1 The Pressure Vessel Design Problem
$g_{1}(\vec{x})=-x_{1}+0.0193x_{3}\leq 0$
$g_{2}(\vec{x})=-x_{3}+0.000954x_{3}\leq 0$
$g_{3}(\vec{x})=-\pi x_{3}^{2}x_{4}+\frac{4}{3}\pi x_{3}^{3}+1296000\leq 0$
$g_{4}(\vec{x})=x_{4}-240\leq 0$
@@ 5.2 Tubular Column Design Problem
$g_{1}(\vec{x})=\frac{P}{\pi x_{1}x_{2}\chi_{y}}-1\leq 0$
$g_{2}(\vec{x})=\frac{8PL^{2}}{\pi^{3}Ex_{1}x_{2}(x_{1}^{2}+x_{2}^{2})}-1\leq 0$
$g_{3}(\vec{x})=\frac{2}{x_{1}}-1\leq 0$
$g_{4}(\vec{x})=\frac{x_{1}}{14}-1\leq 0$
$g_{5}(\vec{x})=\frac{0.2}{x_{2}}-1\leq 0$
$g_{6}(\vec{x})=\frac{x_{2}}{8}-1\leq 0$
@@
6 Model-based Optimization for Real-time Problems
@@
$h_{\theta}(x)$ $=\frac{1}{1+\exp{(-\theta^{T}x)}}$
$J(\theta)$ $=-\frac{1}{m}\sum_{i=1}^{m}[y^{(i)}\ln{(h_{\theta}(x^{(i)}))}+(1-y^{(i)})\ln{(1-h_{\theta}(x^{(i)}))}]$
$\frac{\partial J(\theta)}{\partial\theta_{j}}$ $=\frac{1}{m}\sum_{i=1}^{m}(h_{\theta}(x^{(i)})-y^{(i)})x_{j}^{(i)}$ (20)
@@
$f(\vec{x})=\sum\frac{((\beta_{1}x_{1}+\beta_{2}x_{2})-y)^{2}}{2}$
@@
Where: - $TP$ represents the number of true positives. - $TN$ represents the number of true negatives. - $FP$ represents the number of false positives. - $FN$ represents the number of false negatives.
@@ 7 Conclusion
Overall, the HEO has demonstrated its stable effectiveness and adaptability in addressing general optimization challenges. Future work could focus on further refining the algorithm’s parameters and exploring its application in other practical optimization problems like path planning[36], circuit planning[31]. Additionally, the algorithm could be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications, provide the source code of HEO in: https://github.com/Spike8086/HEO.git, provide the data of the result and the notebook of the experiment in: https://www.kaggle.com/code/spike8086/heo-formal-ver

### v7

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a quantum-inspired metaheuristic designed to address general optimization problems. The HEO mimics the effects between quantum such as tunneling, entanglement. After the introduction to the HEO mechansims, the study presents a comprehensive evaluation of HEO’s performance against extensively-used optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating general optimization problems. The test of HEO in Pressure Vessel Design and Tubular Column Design also infers its feasibility and potential in real-time applications. Further validation of HEO in Osmancik-97 and Cammeo Rice Classification achieves a higher accuracy record.
@@ 1 Introduction
The main motivation behind the development of the Halfway Escape Optimization (HEO) algorithm stems from the limitations of existing optimization methods in terms of efficiency and adaptability to various industrial single-objective optimization problems.
HEO is a quantum-inspired metaheuristic designed to tackle the distinction between different search spaces in industrial scenarios with a focus on achieving fast convergence rates and reducing the time cost in searching. Inspired by the behavior of quantum particles and the concept of halfway escape, the algorithm offers a versatile and adaptive approach to exploration and exploitation in challenging optimization domains.
This paper provides a comprehensive analysis to the HEO algorithm, focusing on its core principles, adaptability, and robustness in navigating diverse optimization landscapes. The unique energy-driven behavior, vibration strategies, and exploratory mechanisms of HEO are examined in the context of solving a range of simple as well as complex benchmark functions. The evaluations demonstrate HEO’s effectiveness in achieving convergence, balancing exploration and exploitation, and discovering high-quality solutions across a variety of challenging optimization problems.
@@ 2 Related Work
There still have other meta-heuristic algorithms, those algorithms could classified as two category: bio-inspired and the nature-inspired, nature-inspired means the optimization algorithms are inspired from physical or chemical laws in nature. Bio-inspired optimizations algorithms including Whale Optimization Algorithm(WOA)[23], Artificial Bee Colony(ABC)[19], Bat Algorithm(BA)[43], Cuckoo Search(CS)[45],Dragonfly Algorithm (DA)[22], Flower Pollination Algorithm(FPA)[44], Cheetah Optimizer (CO)[3], Genetic Programming (GP)[41], Fox Optimizer(FOX)[26], GOOSE algorithm[15].
Nature-inspired optimization algorithms contains Snow Ablation Optimizer (SAO)[8],Light Spectrum Optimization (LSO)[2], Nuclear Reaction Optimization (NRO)[40],Water Cycle Algorithm (WCA)[13],Chernobyl Disaster Optimizer (CDO)[33]. Those bio-inspired as well as nature-inspired algorithms both have contributed significantly to the field of optimization, each with its unique strengths.
@@ 3 Proposed Methodology
In recent years, swarm optimization algorithms have been widely applied in various fields, such as machine learning, engineering design. These algorithms have achieved remarkable success in solving optimization problems and have shown great potential in handling complex tasks.
However, these algorithms often face difficulties when faced with complex landscapes. One example of a group optimization algorithm facing a rugged landscape is the Particle Swarm Optimization (PSO) algorithm. When facing rugged landscapes with many local optima, PSO often gets trapped in these local optima and fails to find the global optimum[38].
@@
Yet QPSO and HEO are both quantum-inspired algorithms with the same mechanism in computing global weight for updating position(seen in equation (6)), they still have significant differences. In contrast to the QPSO, QPSO only allows for a single state change, HEO is more inclined to escape when convergence is achieved, resulting in the inclusion of velocity during the escape phase. Moreover, HEO incorporates random searches not only within the range of updates but also for the entire swarm. Additionally, utilizing a center clipping strategy in HEO contributes to its superior convergence rate.
The Position Update mechanism, which forms the basis of all swarm optimization algorithms, which are the moving way of each individual entity in the swarm, is designed differently in HEO. Instead of focusing on a single peak, the quantum in HEO has two states that determine whether it should approximate the optimum. This unique way of updating positions allows the swarm in HEO to escape from false peaks, thus addressing the first difficulty mentioned earlier. Like the quantum could shuttle through the potential energy barrier [29], the movement of the quantum in HEO is not influenced by gradients like in the PSO algorithm, making it more stable and less likely to be hindered by complex landscapes. Furthermore, in order to navigate complex hollow landscapes, HEO incorporates mechanisms such as increasing the escape factor, Vibration, and Random Skip to accelerate the quantum’s ability to overcome obstacles. Lastly, for exploring complex situations, the Center Clipping mechanism in HEO ensures that the swarm is relatively concentrated around the already discovered optima, thereby efficiently searching for solutions in a convex environment.
@@ 3.2 Position Update
$\mathbb{x_{i+1}}=\mathbb{x_{i}}+\mathbb{v_{g}}+\mathbb{v_{l}},\quad i=1,2,3...N$ (1)
$\mathbb{v_{g}}=(\mathbb{x_{g}}-\mathbb{x_{i}}(c_{i}+1)\times r_{1})\times r_{2}\times r_{3}$ (2)
$\mathbb{v_{l}}=(\mathbb{x_{l}}-\mathbb{x_{i}}(c_{i}+1)\times r_{1})\times r_{2}(1-r_{3})$ (3)
@@ 3.3 Vibration
Another behavior of HEO is common in other algorithms, including AFSA, when the search is meaningless to make a random step for an individual search. Yet those algorithms adopt the strategy, they ignore the influence of size of steps, and those search approaches highly rely on the hyper-parameters tunning. HEO tries to use the standard deviation of positions to judge the magnitude of one random step, creating a relatively stable performance compared to the others, this random search is called Vibration(seen in equation(8).
$\mathbb{x_{i+1}}=\mathbb{x_{i}}+\frac{\mathbb{n}}{1+e^{a_{i}}}$ (8)
@@
$\mathbb{n}\sim N(0,\sigma_{x_{i}})$ (11)
Just like the property of quantum in physics, with higher energy level $a_{i}$, the quantum would be harder get influenced by others[35], which means smaller oscillations(seen in equation (8) and Fig.3), maintain an extent of diversity in the population. To maintain the convergence of the whole group, the increment of energy level $a_{i}$ only works in a few quantum. As for amplitude of vibration, other scaling approaches to adapting the space of high dimensions with gamma functions, are computed much slower than just simply calculating the standard deviation of the position vectors as shown in equation(11). The reason for using Normal distribution due its universal applicability. According to the central limit theorem, with the increasing of the dimension, the size of the vibration is more approximate to the size of the step in the real situation. Since the HEO does not include periodical functions in this behavior, it cannot be defined as oscillations, that’s the reason for its name.
@@ 3.4 Center Clipping
As Fig.4 illustrated, the quantum not only would be restricted by the search area or constraints $S_{bound}$(the area with (a)) with half of the side length $b$, but the group also would be restricted by $S_{group}$. $S_{group}$(the area of (b) in Fig.3)is the search space of the group that HEO allows to search with half of the side length $b_{g}$, the search space in HEO are all in the shape of a square, making it computing faster than calculating sphere. Moreover, as equation (13) shows, this bound of group $S_{group}$ depends on the differences between the current quantum’s position and the best one that HEO found. This center clipping mechanism uses a random factor $r_{5}$, which makes this bound soft edge not converge too early and maintains the diversity of the group. The final result of the group gets into the range of $S_{result}$(area (c)), which is the overlap area of the $S_{group}$(area (b)) and $S_{bound}$(area (a)).
Figure 4: Center Clipping in Ackley Function
@@ 3.5 Random Skip
$\mathbb{x_{i}}=\frac{\mathbb{x_{i}}+\mathbb{\vec{r}}}{2}$ (15)
$\vec{r}_{j}\in\mathbb{\vec{r}}$ (16)
$\vec{r}_{j}\sim U(0,\mathbb{b})$ (17)
The random skip behavior in HEO allows the entire group of quantum to explore a completely new and random search space, the blue area represents the search space, and the orange part is for the constraint. This is achieved by randomly shifting the position of each quantum by a random vector $\vec{r}$ in Equation (15). The components of the random vector $\vec{r}_{j}$ are uniformly sampled from the interval $[0,b]$, as shown in equation (17). There are numerous $b$ for $\vec{r}$, hence the shape of the search space also changed with it, also the space restricts the constraints. This random skip mechanism enables HEO to explore different regions of the search space and potentially discover new and better solutions.
@@ 5.1 The Pressure Vessel Design Problem
$g_{1}(\mathbb{\vec{x}})=-x_{1}+0.0193x_{3}\leq 0$
$g_{2}(\mathbb{\vec{x}})=-x_{3}+0.000954x_{3}\leq 0$
$g_{3}(\mathbb{\vec{x}})=-\pi x_{3}^{2}x_{4}+\frac{4}{3}\pi x_{3}^{3}+1296000\leq 0$
$g_{4}(\mathbb{\vec{x}})=x_{4}-240\leq 0$
@@ 5.2 Tubular Column Design Problem
$g_{1}(\mathbb{\vec{x}})=\frac{P}{\pi x_{1}x_{2}\chi_{y}}-1\leq 0$
$g_{2}(\mathbb{\vec{x}})=\frac{8PL^{2}}{\pi^{3}Ex_{1}x_{2}(x_{1}^{2}+x_{2}^{2})}-1\leq 0$
$g_{3}(\mathbb{\vec{x}})=\frac{2}{x_{1}}-1\leq 0$
$g_{4}(\mathbb{\vec{x}})=\frac{x_{1}}{14}-1\leq 0$
$g_{5}(\mathbb{\vec{x}})=\frac{0.2}{x_{2}}-1\leq 0$
$g_{6}(\mathbb{\vec{x}})=\frac{x_{2}}{8}-1\leq 0$
@@
6 Model-based Optimization for Real-time Problem
@@
$h_{\theta}(\mathbb{x})$ $=\frac{1}{1+\exp{(-\mathbb{\theta}^{T}\mathbb{x})}}$
$J(\mathbb{\theta})$ $=-\frac{1}{m}\sum_{i=1}^{m}[\mathbb{y}^{(i)}\ln{(h_{\mathbb{\theta}}(\mathbb{x}^{(i)}))}+(1-\mathbb{y}^{(i)})\ln{(1-h_{\mathbb{\theta}}(\mathbb{x}^{(i)}))}]$
$\frac{\partial J(\mathbb{\theta})}{\partial\mathbb{\theta}_{j}}$ $=\frac{1}{m}\sum_{i=1}^{m}(h_{\mathbb{\theta}}(\mathbb{x}^{(i)})-\mathbb{y}^{(i)})\mathbb{x}_{j}^{(i)}$ (20)
@@
$f(\mathbb{\vec{x}})=\sum\frac{((\beta_{1}x_{1}+\beta_{2}x_{2})-y)^{2}}{2}$
@@
Where:
- $TP$ represents the number of true positives.
- $TN$ represents the number of true negatives.
- $FP$ represents the number of false positives.
- $FN$ represents the number of false negatives.
@@ 7 Conclusion
Overall, the HEO has demonstrated its stable effectiveness in 14 benchmarks functions and shown adaptability in addressing general optimization like the Pressure Vessel Design Problem, the Tubular Column Design Problem and the Model-based Optimization Problem. Future work could focus on further refining the algorithm’s parameters and exploring its application in other practical optimization problems like path planning[37], circuit planning[32]. The algorithm could also be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications, provide the source code of HEO in: https://github.com/Spike8086/HEO.git